Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-048_2001-054

Mission Operations Days: 2001/048 to 2001/054
February 17, 2001 00:00:00 GMT to February 23, 2001 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)Nominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off; #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68Nominal

This Week’s Completed MODIS Activities:

Saturday, February 17, 2001

None

Sunday, February 18, 2001

None

Monday, February 19, 2001

None

Tuesday, February 20, 2001

None

Wednesday, February 21, 2001

2001/052 21:02:05 ATC OA-16: SDSM Screened
2001/052 22:40:57 ATC OA-15: SDSM Open

Thursday, February 22, 2001

2001/053 20:11:59 ATC OA-27: PV Electronic calibration
2001/053 20:15:29 ATC OA-28: PC Electronic calibration

Friday, February 23, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Saturday, February 24, 2001

None

Sunday, February 25, 2001

None

Monday, February 26, 2001

057/16:15:10 ATC Moved Formatter to Night from 15:44:39 to 16:15:10 to gather approx. 30 minutes of day rate data over groundtrack night

Tuesday, February 27, 2001

None

Wednesday, February 28, 2001

059/21:07:13 ATC OA-16: SDSM Screened 059/22:46:05 ATC OA-15: SDSM Open

Thursday, March 1, 2001

None

Friday, March 2, 2001

None

Maneuvers:

The next MODIS Roll Maneuver for a Lunar Calibration through the Space View Port is scheduled for March 13th at approx. 21:32:52. The next Drag Makeup Maneuver (delta-V) is scheduled for March 14th at approx. 19:40:00

MODIS Anomalies:

None

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Temperature trends are correlating with the solar beta angle.

Non-MODIS Significant Events:

None

Limited Life Item Status:

SRCA 10W Lamp #1: 181.7 of 500 hours
SRCA 10W Lamp #2: 140.5 of 500 hours
SRCA 10W Lamp #3: 149.7 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 556.7 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1537 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements